OpenStudy (anonymous):

use the discriminant to determine the nature of the roots of 8x^2=0 a: no real roots c: one real root b: two distinct real roots d: three distinct real roots

OpenStudy (raden):

D = b^2 - 4ac with D = discriminant

OpenStudy (anonymous):

I don't know how to plug it in though, im horrible at math

OpenStudy (raden):

8x^2=0 here, a = 8, b = 0, and c = 0 so, D = b^2 - 4ac = 0^2 - 4(8)(0) = 0 - 0 = 0 because the discriminant (D) = 0, then it is one real root

OpenStudy (anonymous):

oh ok! ok so I have another one -2x^2-5x-2=0 D= b^2-4ac= 0^2 -4(-2)(0)= 0-0=0????

OpenStudy (anonymous):

Or do I put for b -5 and for c I put -2?

OpenStudy (raden):

the general form of quadratic equation is ax^2 + bx + c = 0 so, if given -2x^2-5x-2=0 it means a = -2, b = -5, and c = -2

OpenStudy (raden):

yeah, a = -2, b = -5, and c = -2 now plug in into formula of D D = b^2 - 4ac

OpenStudy (anonymous):

So D=-5^2 -4(-2)(-2)

OpenStudy (raden):

yeah, D=(-5)^2 -4(-2)(-2) D = 25 - 16 D = 9 it means D > 0, right ?

OpenStudy (anonymous):

I can get the solving part after awhile but what I have trouble with is how do I tell how many roots an equation has

OpenStudy (raden):

there are 3 kinds of roots of quadratic equation : 1) if D > 0 then it has 2 distinct real roots 2) if D = 0 then it has one real root (for example see in your question before) 3) if D < 0 then it has no real roots (or you can call it has 2 imaginary roots)

OpenStudy (anonymous):

Ok dumb question but how do I know if its <, =, >?

OpenStudy (raden):

and dont think this would has three distinct real roots, because the quadratic equation only has maximum 2 roots see, if D = 0 obvious it is the first kind, lol D = 9, 16, 25, 48, 127, 999, or all positive numbers (it is D >0) D = -2, -33, -25, or all negative numbers ( it is D <0)

OpenStudy (raden):

in other words, < 0 actually negative number and > 0 it is positive number

OpenStudy (anonymous):

Got it!!! Thank you very much!!! Wow!! Thanks again!

OpenStudy (raden):

does that helps ?

OpenStudy (raden):

you're welcome :)